OpenStudy (anonymous):

Write with positive exponents 125^-5/3

OpenStudy (anonymous):

When dealing with negatives, take the inverse of the base number to change the negative exponent to a positive. Try it :)

OpenStudy (anonymous):

but i'm not getting the right answer!!! :/

OpenStudy (nubeer):

hmm what is the right answer for it?

OpenStudy (nubeer):

or what your answer is.

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

Alright look, when you have a negative exponent, your goal is to make it positive. You need to take an inverse of the base to do so. What is the inverse of 125?

OpenStudy (anonymous):

-125

OpenStudy (anonymous):

Nope.

OpenStudy (anonymous):

What is 125 in fraction form?

OpenStudy (anonymous):

ohhh sorry 5

OpenStudy (anonymous):

or idk lol i have a serious problem

OpenStudy (anonymous):

Not even that. Alright then lemme teach you.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

The fraction form of any whole number is always \[\frac{ n }{ 1 }\] where n is the number over 1

OpenStudy (anonymous):

So the fraction form of 9 is \[\frac{ 9 }{ 1 }\]

OpenStudy (anonymous):

Do you know what the inverse of a number is?

OpenStudy (anonymous):

isn't it the opposite of the number?

OpenStudy (anonymous):

What is the inverse of 1/4?

OpenStudy (anonymous):

-4 or just 4!!!

OpenStudy (anonymous):

I guess

OpenStudy (anonymous):

4 is the correct answer. The inverse of 1/4 cannot be 4 simply because there is no negative here.

OpenStudy (anonymous):

I mean cannot be -4*

OpenStudy (anonymous):

ohhh okay

OpenStudy (anonymous):

Now, onto your question.

OpenStudy (nubeer):

inverse is basically flipping.. taking numerator in denominator and denominator in numerator.

OpenStudy (anonymous):

What is the inverse of 125?

OpenStudy (anonymous):

1/125

OpenStudy (anonymous):

Good job :) Now that you took the inverse of the base, what happens to the exponent?

OpenStudy (anonymous):

-3/5

OpenStudy (anonymous):

Nope. Remember what I said above. "When dealing with negatives, take the inverse of the base number to change the negative exponent to a positive."

OpenStudy (anonymous):

okayyy so it's 3/5

OpenStudy (anonymous):

Nope. What is the original fraction. You do not need to take the inverse of the exponent. Just the base.

OpenStudy (anonymous):

can you just solve it and when you do i'll look at the process and get it! :/

OpenStudy (anonymous):

Ah, just what I've been saying to the Moderators. Take the inverse: (1/125)^(5/3) Exponent becomes positive. This is because (1/125) = 125^-1 Using the power law for exponents (1/125)^(-1)(-5/3) = (1/125)^(5/3)

OpenStudy (anonymous):

Get it?

OpenStudy (anonymous):

okayyy yes I get it now Thanks and sorry for being slow

OpenStudy (anonymous):

wanna try a practice question?

OpenStudy (anonymous):

wait i have a question! so is it the same for 125^(-5/3)

OpenStudy (anonymous):

try it on your calculator. Remember to put brackets around the exponents or fractions.

OpenStudy (anonymous):

got it?